# A Survey on Monocular Re-Localization: From the Perspective of Scene Map Representation

### v1

@@ Abstract
Monocular Re-Localization (MRL) is a critical component in numerous autonomous applications, which estimates 6 degree-of-freedom poses with regards to the scene map based on a single monocular image. In recent decades, significant progress has been made in the development of MRL techniques. Numerous landmark algorithms have accomplished extraordinary success in terms of localization accuracy and robustness against visual interference. In MRL research, scene maps are represented in various forms, and they determine how MRL methods work and even how MRL methods perform. However, to the best of our knowledge, existing surveys do not provide systematic reviews of MRL from the respective of map. This survey fills the gap by comprehensively reviewing MRL methods employing monocular cameras as main sensors, promoting further research. 1) We commence by delving into the problem definition of MRL and exploring current challenges, while also comparing ours with with previous published surveys. 2) MRL methods are then categorized into five classes according to the representation forms of utilized map, i.e., geo-tagged frames, visual landmarks, point clouds, and vectorized semantic map, and we review the milestone MRL works of each category. 3) To quantitatively and fairly compare MRL methods with various map, we also review some public datasets and provide the performances of some typical MRL methods. The strengths and weakness of different types of MRL methods are analyzed. 4) We finally introduce some topics of interest in this field and give personal opinions. This survey can serve as a valuable referenced materials for newcomers and researchers interested in MRL, and a continuously updated summary of this survey, including reviewed papers and datasets, is publicly available to the community at:
https://github.com/jinyummiao/map-in-mono-reloc
keywords:
Simultaneous Localization and Mapping, Positioning, Monocular Re-Localization, Pose Estimation
@@
1 Introduction
Visual Re-Localization, as a typical state estimation problem, is a challenging and ongoing topic. Given a low-cost monocular camera, Monocular visual Re-Localization (MRL) tends to estimate 6 Degree-of-Freedom (DOF) poses, including orientation and position, with regards to scene map when the vehicle re-visits a mapped area VBL2018PR ; LCD2022TITS . This task holds great attentions in various autonomous applications, e.g., Virtual Reality (VR) Sven2014ECCV ; Jonathan2014TVCG , robots navigation Lim2015IJRR ; LATITUDE ; LocNeRF , Autonomous Driving (AD) TM3Loc ; MLVHM ; DA4AD ; RoadMap . Generally, MRL is commonly discussed to solve long-term association (so called loop closure) problem in Simultaneous Localization and Mapping (SLAM) system LCD2022TITS or recover pose as a state re-initialization in kidnapped robot scenarios Sotirios2020CVC .
As the spatial model of the explored scene, map saves prior knowledge of scene (such as appearance, geometry, topology, etc.) and serves a reference coordinate for MRL. Generally, the MRL methods need to align current sensor data to scene map so that the ego pose can be solved, which can be consistently summarized as a two-step map-matching pipeline: build scene map in a specific form (mapping stage) and then localize based on the matching information between current data and the map (localization stage). Under the context of MRL, the data in the map is captured by sensors in the historical period, and it probably has extreme differences in visual conditions (like weather seqslam , illumination vprbench , day-night Asha2019ICRA , etc.) and even modality mono_i2p_vo ; cmrnet with the currently captured data, challenging the MRL solutions. Therefore, the representation form of data in scene map plays a vital role in MRL task and it affects the robustness of MRL methods against visual interference from real-world.
Figure 2: The structure of this survey.
Under such a context, various kinds of map have been involved in the existing MRL researches to achieve robust, efficient, and high-precision localization. In some proposals aiming at light-weight map, map is represented by frames with geographic location annotation (named as geo-tagged frame map) so that localization can be performed by image retrieval netvlad ; CoHOG or relative pose estimation arnold2022mapfree . For most of MRL methods tending to localization accuracy, map is built by visual point clouds with high-dimensional descriptors (named as visual landmark map) hloc1 ; hloc2 . In this formulation, the alignment can be performed by pixel-wise matching between local features orb ; superpoint . Some cross-modal MRL methods adopt raw point cloud map cmrnet ; mono_i2p_vo for its higher geometry accuracy and density. In the area of AD, the compact and vectorized semantic map, called vectorized High-Definition (HD) map, is widely utilized TM3Loc ; MLVHM . So the alignment process becomes matching between semantic instances. To achieve improvement by the benefit of End-to-End (E2E) way as other computer vision task alexnet ; faster-rcnn , some proposals learn to extract pose-related information from implicit map and localize itself, which takes current image as input and estimates poses posenet ; neumap . We believe that the scene map is implicitly represented in a form of network parameters and name such a kind of map as learnt implicit map. In this study, we find that MRL methods with the same kind of scene map usually perform in a theoretically similar way. Map determines the implementation way and even the localization performance of MRL methods.
However, existing surveys in related field do not discuss about the relationship between MRL methods and the scene map, which blocking the analysis and development of MRL. In the light of such a status, we aim to reduce the gap by reviewing existing MRL methods based on their utilized scene map. Specifically, as shown in Fig. 2, the MRL methods are categorized and introduced in five classes based on its used map, including not only traditional geo-tagged frame map (see Sec. 3), visual landmark map (see Sec. 4), and point cloud map (see Sec. 5), but also vectorized HD Map (see Sec. 6) and recently raised learnt implicit map (see Sec. 7). For each class, we fully review existing methods and deeply analyze the strength, weakness, evaluation metric, public benchmark. To boost the development of the community, we also discuss some opening problem in this field and provide some personal opinions.
  • 1.
  • 2.
    The survey reviews MRL methods from a new perspective that we categorize the existing algorithms based on the representation form of utilized map. The relationships between map and MRL solution can be clearly studied.
  • 3.
    For each kind of MRL method with diverse scene map, we summarize the typical algorithms but also introduce the popularly used datasets, and evaluation metrics.
  • 4.
  • 5.
    As a final contribution, we build a continuously updated repository at github.com, including papers and datasets reviewed in this survey.
Organization The structure of this survey is as follows. Section 1 summarizes the contributions and paper structures. Section 2 describes the research background of MRL. From Sec. 3 to 6, the existing MRL methods are fully reviewed based on their scene map representation. The typical algorithms with various scene map will be evaluated in Sec. 8.2 and we analyze their advantages and disadvantages, respectively. Finally, Section 9 discusses some heated questions about MRL and Section 10 ends with conclusion and promising future directions.
Figure 3: Typical challenges of MRL methods. a) day-night changes with varying illumination tokyo247 , b) seasonal changes with different weather (sunny in summer and snowy in winter) seqslam , c) dramatic viewpoint changes university1652 , d) distinct places with visually similar objects, called perceptual aliasing problem malaga .
2 Background
2.1 Problem Formulation and Symbols Definition
We start by define MRL problem in theory and introduce the symbols used in this survey.
During the offline mapping stage, we build the scene map in a specific representation form. In those MRL method that scene map is explicitly represented and saved, scene map is $\mathcal{M}^{\mathcal{G}}=\left\{\textbf{m}^{\mathcal{G}}_{i}\right\}$, where $\textbf{m}^{\mathcal{G}}_{i}$ is the $i$-th represented item in the map, e.g., visual landmarks, point clouds, and semantics. Here, {}$\{\cdot\}^{\mathcal{G}}$ denotes the map is considered in a global coordinate of this scene $\mathcal{G}$ by default and can be simplified as =$\mathcal{M}=\{\textbf{m}_{i}\}$ if not specified. When a vehicle $v$ revisits such a mapped area, it captures the observation of the scene, that is, a monocular image $\mathcal{I}$ in MRL research. The vehicle $v$ can estimate its current ego-pose based on the current observations $\mathcal{I}$ given scene map information $\mathcal{M}$, which is defined as a MRL problem. Formally, the MRL problem can be written as
$\hat{\textbf{x}}^{\mathcal{G}}_{v}=\mbox{MRL}(\mathcal{I}~{}|~{}\mathcal{M})$ (1)
where ${}^^\hat{\{\cdot\}}$ denotes an estimated value, $\mbox{MRL}(\cdot)$ is a MRL solution, and $\textbf{x}^{\mathcal{G}}_{v}$ is the vehicle’s ego-pose with regard to $\mathcal{G}$.
Commonly, a 6 DOF pose $\textbf{x}^{\mathcal{G}}_{v}=\left[\textbf{R}^{\mathcal{G}}_{v}~{};~{}\textbf{t}^{\mathcal{G}}_{v}\right]$ is composed by a 3 DOF rotation $\textbf{R}^{\mathcal{G}}_{v}$ and a 3 DOF translation $\textbf{t}^{\mathcal{G}}_{v}$. The rotation can be represented as quaternion, rotation matrix, or Euler Angle. In some MRL application for ground AD vehicles, only 3 DOF pose is considered, i.e., lateral and longitudinal position $[x,y]$ and heading/yaw angle $\theta_{yaw}$.
For better readability, we define frequently used symbols here. The current image in localization stage (query image) is defined as $\mathcal{I}_{q}$ while a historical image at $t$ timestamp in mapping stage (reference image) is defined as $\mathcal{I}_{r(t)}$. Local features from $\mathcal{I}_{t}$ are denoted as $\mathcal{F}^{t}=\{\textbf{f}^{t}_{i}\}$ where a local feature $\textbf{f}^{t}_{i}$ contains a key point $\textbf{p}^{t}_{i}=[u,v]\in\mathbb{R}^{2}$ and corresponding local descriptor $\textbf{d}^{t}_{i}\in\mathbb{R}^{C}$ where $C$ is the dimension. A pair of matched local features between $\mathcal{I}_{q}$ and $\mathcal{I}_{r(t)}$ is denoted as $<\textbf{f}^{q}_{i},\textbf{f}^{r(t)}_{j}>$. The global feature of $\mathcal{I}_{t}$ is a global descriptor and denoted as $\mathcal{F}^{t}\in\mathbb{R}^{C}$. The items in $\mathcal{M}$ share the same symbols but in global coordinate, i.e., {}$\{\cdot\}^{\mathcal{G}}$.
2.2 Challenges of Monocular Re-Localization
  • 1.
    Appearance change: Since monocular camera serves as the primary sensor in MRL solutions, the appearance changes of scene will dramatically affect the algorithms. The conditional appearance variance, e.g., weather seqslam , illumination vprbench ; tokyo247 , seasons seqslam ; 4seasons , and day-night changes benchmark6dof makes current image visually different from the map, thus challenging matching. And structural changes like dynamic occlusions slcd , and layout changes InLoc will interfere with the geometrical pose estimation.
  • 2.
    Viewpoint difference: When the camera viewpoint is quite different between current timestamp and mapping stage, MRL method will be struggled by limited co-visible area and changed layout of items. As typical instances, cross-view MRL for drones each_part_matters ; university1652 , opposite-view donot_look_back ; garg2018lost and even ground-to-aerial xview MRL for ground vehicles attract growing interests.
  • 3.
    Perceptual aliasing: In some scenaoris with visually similar or repeated textures, the MRL method will generate ambiguous estimation when distinct places have similar appearances, e.g., corridors rapnet ; openloris , and parking lots malaga , which is called perceptual aliasing problem.
  • 4.
    Generalization and scalability: The real world scenarios are infinite, and we cannot exhaust all the types, visual conditions, and interference of possibily occurred scenes. A practical real-world autonomy requires MRL methods to work stably in diverse scenarios, even in unseen environments, called generalization ability keetha2023anyloc ; SANet . Besides, the real world is unbounded, we also need MRL solutions scalable, which limits the unbearable increasement of map size and computational costs with vehicles continuously explore.
2.3 Comparison with other surveys
In decades of years, numerous reviews on autonomous robotics techniques have been raised, addressing various topics, e.g., place recognition LCD2022TITS ; Yin2022GeneralPR ; VPR_a_survey ; vpr_survey_ijcai ; VRP_a_survey_from_DL_perspective , ego and object localization VBL2018PR ; a_survey_on_visual_map_localization_using_LiDARS_and_camera ; visual_and_object_geo_localization_a_comprehensive_survey , SLAM past_present_future_SLAM ; a_survey_on_DL_for_SLAM ; visual_SLAM . As a milestone survey, Lowry et al. VPR_a_survey firstly introduced the concepts behind visual place recognition. Later, Garg et al. vpr_survey_ijcai refine the definition by introducing orientation limitation. With a growing number of visual place recognition methods based on deep learning, Zhang et al. VRP_a_survey_from_DL_perspective reviewed recently proposed solutions from the deep learning perspective. Yin et al. Yin2022GeneralPR presented general place recognition challenges. However, place recognition methods could only serve as a coarse localization that provide a rough pose approximation, which only cover few part of visual localization researches. In a_survey_on_visual_map_localization_using_LiDARS_and_camera , Elhousni et al. defined visual map localization as a two-stage process, namely, place recognition and metric map localization, and reviewed MRL methods using LiDAR, camera, or cross-modal sensors. Such a two-stage framework is only adopted by those MRL solutions using visual landmark map and point cloud map, and it cannot completely describe all the MRL solutions, such as newly emerged absolute pose regression-based methods posenet , scene coordinate regression-based methods neumap , etc.
By deeply analyze existing reviews about MRL methods, we find that although all these surveys have made great contribution to the community and boosted the progress of MRL researches, they have regrettably neglected the relationship between MRL methods and scene map, and thus cannot analyze MRL solutions in a unified framework. In this paper, we propose to review MRL methods from the perspective of scene map so that we can clearly categorize them based on the representation form of utilized map, thereby providing a comprehensive and deep review about MRL methods.
3 Geo-tagged Frame Map
During mapping stage, many frames are labeled by geographic position or even precise 6 DoF poses. These geo-tagged frames can be served as scene map in MRL methods. In this field, some proposals estimate ego-pose by historical pose of retrieved geo-tagged frames (Fig. 4 a) or estimated relative pose between query image and geo-tagged frame (Fig. 4 b).
3.1 Visual Place Recognition
Visual Place Recognition (VPR) algorithm aims to identify the re-observed places by retrieving reference frame when the vehicle goes back to a previously visited scene vpr_survey_ijcai ; VPR_a_survey , which is also adopted for loop closure detection in SLAM system LCD2022TITS . VPR regards the pose of retrieved reference frame as an initial or approximated pose of current query image. Formally, in VPR, the scene map $\mathcal{M}$ is represented by reference frames $\mathcal{I}_{r(t)}$ with pose $\textbf{x}^{\mathcal{G}}_{t}$, i.e., =$\mathcal{M}=\{\textbf{m}_{t}\}$ where $\textbf{m}_{t}=<\mathcal{I}_{r(t)},\textbf{x}^{\mathcal{G}}_{t}>$. Generally, given the query image $\mathcal{I}_{q}$, a VPR algorithm tends to retrieve the best-matched reference frame based on the image similarity:
$\hat{\textbf{m}}=\mathop{\arg\max}\limits_{\mathbf{m}}P(\mathcal{I}_{q},\mathcal{I}_{r(i)}~{}\mbox{of}~{}\textbf{m}_{i}|\textbf{m}_{i}\in\mathcal{M})$ (2)
@@ 3.1 Visual Place Recognition
where the matching function $P(A,B)$ provides the similarity or matching score between $A$ and $B$, and $\textbf{m}_{i}$ is a matching candidate, $\hat{\textbf{m}}$ is the best-matched one for $\mathcal{I}_{q}$. Then, the pose of current query image can be given by the pose of retrieved reference frame:
@@
In early proposals, SeqSLAM seqslam and Fast-SeqSLAM fastslam direcly use down-sampled and normalized image patches to measure image similarity. However, raw image intensity is sensitive to visual interference. So VPR methods often extract high-level image features from images and measure feature similarity as a more reliable image similarity. In this feature-based scheme, image feature algorithm plays an important role in VPR solutions. In VPR, based on the receptive field, the image feature can be categorized into global feature hog ; vlad ; netvlad and local feature sift ; superpoint .
Global Feature-based VPR (GF-VPR): Global feature algorithm can directly represent the whole image as a compact matrix or vector so that the image similarity can be simply defined as a cosine distance $P(A,B)=({A\cdot B})/({\|{A}\|\|{B}\|})$ in Eqn. 2. Therefore, the core component of the GF-VPR methods should be the global feature algorithm. Before the deep learning stage, the global image features utilized in this area is designed to analyze the statistical distribution of image intensity, e.g., Gist Siagian , color histograms Ulrich , HOG hog . Zaffer et al. CoHOG improved HOG-based GF-VPR method by using image entropy to extract Regions-of-Interest and regional-convolutional descriptor matching. As a widely used solution, VLAD vlad converts $N$ local descriptors $\{\textbf{d}_{i}\}^{N}_{i=1}$ into a compact matrix $V$ using a codebook with $K$ clusters:
$V(j,k)=\sum^{N}_{i=1}a_{k}(x_{i})\left(\textbf{d}_{i}(j)-\textbf{c}_{k}(j)\right)$ (4)
where $\textbf{d}_{i}(j)$ and $\textbf{c}_{k}(j)$ are the $j$-th dimensions of the $i$-th descriptor and $k$-th cluster centre, respectively. $a_{k}(x_{i})$ denotes the membership of the $\textbf{d}_{i}$ to $\textbf{c}_{k}$, i.e., $a_{k}(\textbf{d}_{i})=1$ if $\textbf{c}_{k}$ is the closest cluster to $\textbf{d}_{i}$ and 0 otherwise.
Since deep learning-based methods dominate the visual tasks, some VPR methods have applied Deep Neural Network (DNN) to extract global features. As a milestone of DNN-based global feature, NetVLAD netvlad designs a trainable feature aggregation layer based on the principle of VLAD algorithm vlad . It replaces the non-differentiable VLAD pooling with soft-assignment of descriptors to multiple clusters, and construct the NetVLAD layer amenable to E2E training via back-propagation:

### v3

@@ Abstract
Monocular Re-Localization (MRL) is a critical component in autonomous applications, estimating 6 degree-of-freedom ego poses w.r.t. the scene map based on monocular images. In recent decades, significant progress has been made in the development of MRL techniques. Numerous algorithms have accomplished extraordinary success in terms of localization accuracy and robustness. In MRL, scene maps are represented in various forms, and they determine how MRL methods work and how MRL methods perform. However, to the best of our knowledge, existing surveys do not provide systematic reviews about the relationship between MRL solutions and their used scene map representation. This survey fills the gap by comprehensively reviewing MRL methods from such a perspective, promoting further research. 1) We commence by delving into the problem definition of MRL, exploring current challenges, and comparing ours with existing surveys. 2) Many well-known MRL methods are categorized and reviewed into five classes according to the representation forms of utilized map, i.e., geo-tagged frames, visual landmarks, point clouds, vectorized semantic map, and neural network-based map. 3) To quantitatively and fairly compare MRL methods with various map, we introduce some public datasets and provide the performances of some state-of-the-art MRL methods. The strengths and weakness of MRL methods with different map are analyzed. 4) We finally introduce some topics of interest in this field and give personal opinions. This survey can serve as a valuable referenced materials for MRL, and a continuously updated summary of this survey is publicly available to the community at:
github.com/jinyummiao/map-in-mono-reloc
Index Terms:
Simultaneous Localization and Mapping, Scene Map, Monocular Re-Localization, Pose Estimation
I Introduction
Visual Re-Localization, as a typical state estimation problem, is a challenging and ongoing topic. Using a low-cost monocular camera, Monocular Re-Localization (MRL) tends to estimate 6 Degree-of-Freedom (DoF) poses, including orientation and position, with regards to scene map when the vehicle re-visits a mapped area [1, 2]. This task holds great attentions in various autonomous applications, e.g., Virtual Reality (VR) [3, 4], robots navigation [5, 6, 7], Autonomous Driving (AD) [8, 9, 10, 11]. Generally, MRL is commonly discussed to solve long-term association (so called loop closure) problem in Simultaneous Localization and Mapping (SLAM) system [2], recover ego pose as a state re-initialization step in kidnapped robot scenarios [12], or retrieve prior information from a pre-built map to complement online perception systems [13].
As the spatial model of the explored scene, map saves prior knowledge of scene (such as appearance, geometry, topology, etc.) and serves a reference coordinate for MRL. Generally, the MRL methods align currently observed sensory data to scene map to solve ego pose, which can be summarized as: build scene map in a specific representation form (mapping stage) and then localize based on the matching information between current observations and the map (localization stage). In MRL solution, the data in the map is captured by sensors in the historical period, and it probably has extreme differences in visual conditions (like weather [14], illumination [15], day-night change [16], etc.) and even modality [17, 18] with the currently observed data, making MRL a technically hard problem. Therefore, the representation form of data in scene map plays a vital role in MRL task and affects the robustness of MRL methods against real-world interference.
Figure 2: The structure of this survey.
Various kinds of map have been involved in the existing MRL researches to achieve robust, efficient, and high-precision localization. In some applications without 3-dimensional (3D) scene model, map is represented by frames with geographic location annotation (named as geo-tagged frame map) so that localization can be achieved by image retrieval [19, 20] or relative pose estimation [21, 22]. For most of MRL methods tending to high localization accuracy, map is composed of visual point clouds with high-dimensional descriptors (named as visual landmark map) [23, 24]. In this formulation, the alignment can be performed by pixel-wise matching between current observations and map data. Some cross-modal MRL methods adopt LiDAR point cloud map [18, 17] for its high geometry accuracy and density. In the area of AD, the compact and vectorized semantic map, called vectorized High-Definition (HD) map, is widely utilized [8, 9]. So the alignment process becomes instance-wise matching between semantic map elements. To achieve improvement by the benefit of End-to-End (E2E) scheme as other computer vision task [25, 26], some methods learn to extract pose-related information from implicit map and localize itself, which no longer require explicit scene map [27, 28]. We believe that the scene maps of these methods are implicitly represented in a form of network parameters and name such a kind of map as learnt implicit map. In this study, we find that MRL methods with the same kind of scene map usually perform in a theoretically similar way: the map representation determines the implementation way and even the localization performance of MRL methods.
However, existing surveys in related field do not discuss about the relationship between MRL methods and the scene map, which blocking the analysis and development of MRL. In the light of such a status, we aim to reduce the gap by reviewing existing MRL methods based on their utilized scene map. Specifically, as shown in Fig. 2, the MRL methods are categorized and introduced in five classes based on its used map, including not only traditional geo-tagged frame map (Sec. III), visual landmark map (Sec. IV), and point cloud map (Sec. V), but also vectorized HD Map (Sec. VI) and recently raised learnt implicit map (Sec. VII). For each kind, we comprehensively review related MRL methods and deeply analyze their strengths, weaknesses, evaluation metrics, and public benchmarks. To further boost the development of the community, we also discuss some opening problems in this field and provide some personal opinions.
  • The survey reviews MRL methods from a new perspective that we categorize the existing algorithms based on the representation form of utilized map. The relationships between map and MRL methods can be clearly studied.
  • For each kind of MRL method with diverse scene map, we summarize the typical algorithms but also introduce the popularly used datasets and evaluation metrics.
  • As a final contribution, we build a continuously updated repository at github.com, including landmark papers reviewed in this survey.
Organization The structure of this survey is as follows. Section I summarizes the contributions and paper structures. Section II describes the research background of MRL problem. From Sec. III to VII, the existing MRL methods are fully reviewed based on their scene map representation. The typical algorithms with various scene map will be evaluated in Sec. VIII and we analyze their advantages and disadvantages, respectively. Finally, Section IX discusses some prominent questions about MRL and section X ends with conclusions.
Figure 3: Typical challenges of MRL methods. a) day-night changes with varying illumination [29], b) seasonal and weather changes (sunny in summer v.s. snowy in winter) [14], c) viewpoint changes [30], d) distinct places with visually similar appearances (perceptual aliasing problem) [31].
II Background
II-A Problem Formulation and Symbols Definition
We start by defining MRL problem in theory and introducing the symbols used in this survey.
During the offline mapping stage, we build the scene map in a specific representation form. In those MRL method that scene map is explicitly represented and saved, scene map can be formatted as $\mathcal{M}_{\mathcal{G}}=\left\{\textbf{m}^{\mathcal{G}}_{i}\right\}$, where $\textbf{m}^{\mathcal{G}}_{i}$ is the $i$-th represented item in the map, e.g., visual landmarks, point clouds, and semantics. Here, {}$\{\cdot\}^{\mathcal{G}}$ denotes the map is considered in a global coordinate $\mathcal{G}$ by default and can be simplified as =$\mathcal{M}=\{\textbf{m}_{i}\}$ if not specified. When a vehicle $v$ revisits such a mapped area, it captures the observation of the scene, that is, a monocular image $\mathcal{I}$ in MRL research. The vehicle $v$ tends to estimate its current ego pose based on the current observations $\mathcal{I}$ given scene map $\mathcal{M}$, which is defined as a MRL problem:
$\hat{\textbf{x}}^{\mathcal{G}}_{v}=\mbox{MRL}(\mathcal{I}~{}|~{}\mathcal{M})$ (1)
@@ II-A Problem Formulation and Symbols Definition
Commonly, a 6 DoF pose $\textbf{x}^{\mathcal{G}}_{v}=\left[\textbf{R}^{\mathcal{G}}_{v};~{}\textbf{t}^{\mathcal{G}}_{v}\right]$ is composed by a 3 DoF rotation $\textbf{R}^{\mathcal{G}}_{v}$ and a 3 DoF translation $\textbf{t}^{\mathcal{G}}_{v}$. The rotation can be represented as quaternion, rotation matrix, or Euler angle. In some MRL application for ground AD vehicles, only 3 DoF pose is considered, i.e., lateral and longitudinal position $[x,y]$ and heading/yaw angle $\theta_{yaw}$.
For better readability, we define frequently used symbols here. The current image in localization stage is defined as query image $\mathcal{I}_{q}$ while a historical image at $t$ timestamp in mapping stage is defined as reference image $\mathcal{I}_{r(t)}$. Local features from $\mathcal{I}_{t}$ are denoted as $\mathcal{F}^{t}=\{\textbf{f}^{t}_{i}\}$ where a local feature $\textbf{f}^{t}_{i}$ contains a key point $\textbf{p}^{t}_{i}=[u,v]\in\mathbb{R}^{2}$ and a corresponding local descriptor $\textbf{d}^{t}_{i}\in\mathbb{R}^{C}$ where $C$ is the dimension. A pair of matched local features between $\mathcal{I}_{q}$ and $\mathcal{I}_{r(t)}$ is denoted as $<\textbf{f}^{q}_{i},\textbf{f}^{r(t)}_{j}>$. The global feature of $\mathcal{I}_{t}$ is a global descriptor and denoted as $\mathcal{F}_{t}\in\mathbb{R}^{C}$. The items in $\mathcal{M}$ share the same symbols but in global coordinate, i.e., {}$\{\cdot\}^{\mathcal{G}}$.
II-B Challenges of Monocular Re-Localization
  • Appearance change: Since monocular camera serves as the primary sensor in MRL solutions, the appearance changes of scene will dramatically affect the algorithms. The conditional appearance variance, e.g., weather [14], illumination [15, 29], seasons [14, 32], and day-night changes [33] makes current image visually different from the map, thus challenging matching. And structural changes like dynamic occlusions [34], and layout changes [35] will interfere with the geometrical pose estimation.
  • Viewpoint difference: When the camera viewpoint is quite different between current timestamp and mapping stage, MRL method will be struggled by limited co-visible area and changed layout of items. As typical instances, cross-view MRL for drones [36, 30], opposite-view [37, 38] and even ground-to-aerial [39] MRL for ground vehicles attract growing interests.
  • Perceptual aliasing: In some scenarios with visually similar or repeated textures, the MRL method will generate ambiguous estimation when distinct places have similar appearances, e.g., corridors [40, 41], and parking lots [31], which is called perceptual aliasing problem.
  • Generalization and scalability: The real world scenarios are infinite, and we cannot exhaust all the types, visual conditions, and interference of possibly occurred scenes. A practical real-world autonomy requires MRL methods to work stably in diverse scenarios, even in unseen environments, called generalization ability [42, 43]. Besides, the real world is unbounded, we also need MRL solutions scalable, which limits the unbearable increasement of map size and computational costs with vehicles continuously explore.
II-C Comparison with other surveys
In decades of years, numerous reviews on autonomous robotics techniques have been raised, addressing various topics, e.g., place recognition [2, 44, 45, 46, 47], ego and object localization [1, 48, 49], SLAM [50, 51, 52]. As a milestone survey, Lowry et al. [45] firstly introduced the concepts behind visual place recognition. Later, Garg et al. [46] refined the definition by introducing orientation limitation. With a growing number of visual place recognition methods based on Deep Learning (DL), Zhang et al. [47] reviewed recently proposed solutions from the DL perspective. Yin et al. [44] presented general place recognition challenges. However, place recognition methods could only serve as a coarse localization that provide a rough pose approximation, which only cover few parts of visual localization researches (Sec. III-A in this survey). Place recognition methods are reviewed and called indirect MRL methods in [1], and the authors also provide direct MRL methods that give the exact 6 DoF pose of the query according to a known reference, which corresponds to Sec. IV, VII-A, and VII-B in this survey. In [48], Elhousni et al. defined visual map localization as a two-stage process, namely, place recognition and metric map localization, and reviewed MRL methods using LiDAR, camera, or cross-modal sensors. Such a two-stage framework is only adopted by those MRL solutions using visual landmark map and point cloud map, and it cannot completely describe all the MRL solutions, such as newly emerged absolute pose regression-based methods [27], scene coordinate regression-based methods [28], etc.
By deeply analyze existing reviews about MRL methods, we find that although all these surveys have made great contributions to the community and boosted the progress of MRL research, they have regrettably neglected the relationship between MRL methods and scene map, and thus cannot analyze MRL solutions in a unified framework. In this paper, we propose to review MRL methods from the perspective of scene map so that we can clearly categorize them based on the representation form of utilized map, thereby providing a comprehensive and deep review about MRL methods.
III Geo-tagged Frame Map
During mapping stage, many frames are labeled by geographic position or precise 6 DoF poses. These geo-tagged frames can be served as scene map in MRL methods. In this field, some proposals estimate ego pose by historical pose of retrieved geo-tagged frames as shown in Fig. 4 a) and others estimate relative pose between query image and geo-tagged frames as shown in Fig. 4 b).
III-A Visual Place Recognition
Visual Place Recognition (VPR) algorithm aims to identify the re-observed places by retrieving reference frame when the vehicle goes back to a previously visited scene, which is also adopted for loop closure detection in SLAM system [2]. After obtaining best-matched reference image, VPR regards the pose of retrieved reference frame as an approximated pose of current query image.
Formally, in VPR, the scene map $\mathcal{M}$ is represented by reference frames $\mathcal{I}_{r(t)}$ with pose $\textbf{x}^{\mathcal{G}}_{t}$, i.e., =$\mathcal{M}=\{\textbf{m}_{t}\}$ where $\textbf{m}_{t}=<\mathcal{I}_{r(t)},\textbf{x}^{\mathcal{G}}_{t}>$. Given query image $\mathcal{I}_{q}$, a VPR algorithm
  • 1
    represent $\mathcal{I}_{q}$ and all the $\mathcal{I}_{r(t)}$ in $\mathcal{M}$ as features:
@@ III-A Visual Place Recognition
  • 2
    retrieve the best-matched reference image based on the similarity of features:
@@
where the matching function $P(A,B)$ provides the matching similarity between $A$ and $B$, and $\textbf{m}_{i}$ is a matching candidate, $\hat{\textbf{m}}$ is the best-matched one for $\mathcal{I}_{q}$. Then, the pose of current query image is given by the pose of retrieved reference frame:
@@
In early proposals, SeqSLAM [14] and Fast-SeqSLAM [53] directly use down-sampled and normalized image patches to measure image similarity. However, raw image intensity is sensitive to visual interference. So VPR methods often extract high-level image features from images and measure feature similarity as a more reliable image similarity. In this feature-based scheme, image feature algorithm plays an important role in VPR solutions. In VPR, based on the receptive field, the image feature can be categorized into global feature [54, 55, 19] and local feature [56, 57].
Figure 5: Four kinds of image features in VPR, including a) image global feature, b) image global feature, c) image sequence feature, and d) image semantics.
Global Feature-based VPR (GF-VPR): Global feature algorithm directly represent the whole image as a compact matrix or vector so the image similarity can be simply measured as a cosine distance
$P(A,B)=({A\cdot B})/({\|{A}\|\|{B}\|})$ (5)
Therefore, the core component of the GF-VPR methods is the feature algorithm $\mbox{ImFeat}(\cdot)$. Before the DL stage, the global image features utilized in this area is designed to analyze the statistical distribution of image intensity, e.g., Gist [58], color histograms [59], VLAD [55], BoW [60], ASMK [61], HOG [54], and improved CoHOG [62]. As a widely used instance, VLAD [55] aggregates $N$ local descriptors $\{\textbf{d}_{i}\}^{N}_{i=1}$ into a compact matrix $V$ using a codebook with $K$ clusters:
$V(j,k)=\sum^{N}_{i=1}a_{k}(\textbf{d}_{i})\left(\textbf{d}_{i}(j)-\textbf{c}_{k}(j)\right)$ (6)
where $\textbf{d}_{i}(j)$ and $\textbf{c}_{k}(j)$ are the $j$-th dimensions of the $i$-th descriptor and $k$-th cluster center, respectively. $a_{k}(\textbf{d}_{i})$ denotes the membership of the $\textbf{d}_{i}$ to $\textbf{c}_{k}$, i.e., $a_{k}(\textbf{d}_{i})=1$ if $\textbf{c}_{k}$ is the closest cluster to $\textbf{d}_{i}$ and 0 otherwise.
Since DL-based methods dominate the visual tasks, some VPR methods have applied Deep Neural Network (DNN) to extract global features. As a milestone of DNN-based global feature, NetVLAD [19] designs a trainable feature aggregation layer based on the principle of VLAD algorithm [55]. It replaces the non-differentiable function $a_{k}(\textbf{d}_{i})$ with soft-assignment of descriptors to multiple clusters, and thus construct an fully-differentiable NetVLAD layer: